# Transverse diffeomorphism invariant spin-2 theories via gauge invariants

### v1

@@ Abstract
We analyze the particle spectrum of a second-order (in derivatives) theory based on a rank-2 tensor field with both symmetric and antisymmetric components. By demanding the existence of a propagating massless spin-2 particle and invariance under transverse diffeomorphisms, we derive a new class of stable models with two massless scalars and a single massless spin-2 particle. A natural nonlinear completion is proposed in terms of a dynamical metric field. The identification of the spectrum is carried out using a fully Lagrangian, gauge-invariant approach which makes use of Bardeen variables in a constructive manner. The approach significantly reduces the number of steps in the spectrum determination in some cases.
@@ 1 Introduction
The TDiff models have some general interesting aspects. Notably, they have an integration constant with the same mass dimension as the cosmological constant which may have interesting phenomenological consequences in scale invariant theories of gravity [8]. In WTDiff gravity, the constant plays the role of a cosmological constant that does not receive quantum corrections, see discussion in [7]. It is also worth commenting that TDiff models (with or without Weyl symmetry) cannot become massive in their spin-2 sector [5]. Given the stringent experimental bounds on the graviton mass ($m_{g}\lesssim 10^{-23}~{}\mathrm{eV}$), the TDiff symmetry seems to be more natural than the usual Diff one (Einstein-Hilbert) which can become massive at linearized level [9] and beyond [10, 11], see also [12].
@@
Here we merge the previous directions of modifying gravity at linearized level and investigate the most general Lorentz invariant quadratic theory of second order in derivatives invariant under TDiff and described by a full rank-2 tensor. The content of the model is examined by using an explicitly gauge invariant approach based on an improved use of Bardeen variables [19].
@@ 2.1 TDiff spectrum via a shortcut
The most general massless TDiff invariant local spin-2 model of second order in derivatives, described only in terms of a symmetric rank-2 field, is given by the following action:
@@
The first two terms are required for massless spin-2 propagation without spin-1 ghosts [5]. The last two terms depend on two real constants: $(a,b)$. Up to surface terms, neglected throughout this work, the action (1) is TDiff invariant444 Here we use the index “$T$” for space-time ($\partial^{\mu}v_{\mu}^{T}$=0) and “$t$” for space ($\partial_{j}v_{j}^{t}=0$) transverse quantities. ($\partial^{\mu}\epsilon^{T}_{\mu}=0$) for arbitrary $(a,b)$,
@@
In the first case, we have the linearized Einstein-Hilbert [^5] More precisely, there is a continuous family of models $(a(r),b(r))$ equivalent to the point $(a,b)=(1,1)$ via what we call r-shifts: $h_{\mu\nu}\to h_{\mu\nu}+r\,\eta_{\mu\nu}\,h$ for which the enhancement still holds, though (3) is modified accordingly, while the case $(a,b)=(a_{D},b_{D})$ is a fixed point of the r-shifts [5]. theory $\mathcal{L}_{TD}(1,1)=\mathcal{L}_{EH}$ where TDiff is enhanced to Diff. In the second one, the symmetry becomes Weyl + TDiff (WTDiff). In both cases, the model describes only massless spin-2 particles in $D=4$ with only two degrees of freedom corresponding to the helicities ±$\pm 2$.
For an otherwise arbitrary couple $(a,b)$, a quite lengthy analysis, see [5],[20] and [21], reveals that we have in general $(D-1)(D-2)/2$ degrees of freedom corresponding to a massless spin-2 and a massless spin-0 particle in $D=4$. A key quantity is $f_{D}(a,b)$:
@@
where $A,B=0,\cdots,N_{I}$. The matrix ${\cal M}_{AB}$ depends on space and time derivatives. The invariant indices $A,B$ are just labeling indices. They are deprived of meaning regarding Lorentz transformations. Regarding their range $N_{I}$, it is the number of independent gauge invariants. We will show that $N_{I}$ equals the number of fields minus the number of independent gauge parameters, generically,
@@
When we go from Diff to TDiff we have one less independent gauge parameter due to $\partial^{\mu}\epsilon_{\mu}^{T}=0$, consequently, from (7) we must have one extra gauge invariant. Since any Diff invariant must be also TDiff invariant and the trace $h$ is TDiff invariant by itself, we can consider $I_{\tilde{A}}^{TDiff}=\left\{I_{A}^{Diff},h\right\}$ as a basis of independent TDiff invariants. The trace $h$ cannot be a function of the Diff invariants $I_{A}^{Diff}$, otherwise it would be also Diff invariant which is certainly not the case, recall $\delta\,h=2\,\partial\cdot\epsilon\neq 0$. So if we were able to rewrite (1) as a sum of the linearized EH theory and a term that depends only on the trace $h$ we could immediately identify the particle content of (1). Indeed, there is a unique field redefinition which accomplishes it, namely, after $h_{\mu\nu}\to h_{\mu\nu}+r\,\eta_{\mu\nu}h$ with $r=(1-a)/(aD-2)$, assuming $a\neq 2/D$, the TDiff theory (1) can be written, in terms of the new field, as
@@
Thus, although the field $h_{\mu\nu}$ in ${\cal L}_{EH}(h_{\mu\nu})$ is not traceless, it turns out that the particle content of ${\cal L}_{TD}(a,b)$ can be read off from the two independent terms in (8), in full agreement with the more laborious analysis of [5], [20] and [21].
@@
where ${\cal L}_{WTD}(h_{\mu\nu})={\cal L}_{TD}(a_{D},b_{D})$, with $f_{D}(a_{D},b)=(D-2)(b_{D}-b)$ and $I_{A}^{WTD}$ are the invariants under WTDiff. Notice that $N_{I}^{WTD}=N_{I}^{Diff}$ since the longitudinal component of $\epsilon_{\mu}$ is replaced by the Weyl gauge parameter such that the number of independent gauge parameters remains the same. On the other hand, as we go from WTDiff to TDiff, the broken Weyl symmetry gives rise to one extra invariant which may be again the trace $h$ that now breaks the Weyl symmetry such that we can choose $I_{\tilde{A}}^{TDiff}=\left\{I_{A}^{WTD},h\right\}$ as another basis for the TDiff invariants. The physical content of ${\cal L}(a_{D},b)$ factorizes into the content of the two terms in (9), once again in agreement with [5], [20] and [21].
@@ 2.2 Diff, WTDiff and TDiff gauge invariants
we can express $\epsilon_{\mu}=\epsilon_{\mu}(\delta h_{\mu\nu})$. It is convenient to decompose the gauge parameters in terms of $SO(D-1)$ scalars and vectors,
@@
where $\omega_{ij}=\partial_{i}\partial_{j}/\nabla^{2}$ and $\theta_{ij}=\delta_{ij}-\omega_{ij}$ are longitudinal and transverse space projection operators. Using (12) into (10), the remaining $D(D+1)/2-D$ equations become invariance equations $\delta I_{A}^{Diff}=0$, where $I_{A}^{Diff}=\{I_{ij}^{t},I_{j}^{t},I_{00}\}$ are the Diff invariants:
@@
They represent a total of $N_{I}^{Diff}=1+(D-2)+[(D-1)D/2-(D-1)]=(D-1)D/2$ independent Diff invariants in agreement with (7): $N(h_{\mu\nu})-N(\epsilon_{\mu})=D(D+1)/2-D$. It is convenient to redefine them on physical grounds. Since the two helicity states ±$\pm 2$ in $D=4$ correspond to the traceless and transverse components of $h_{ij}$, we decompose $I_{ij}^{t}$ into its traceless ($I_{ij}^{tt}$) and trace ($I$) components. Respectively,
@@
$I=(D-2)\nabla^{2}\Lambda\ ;\qquad h=-\rho+\nabla^{2}(D-2)\Lambda+\nabla^{2}\psi$ (21)
We can always associate $I_{ij}^{tt}$ and $I_{j}^{t}$ with the field components $\Psi_{ij}^{tt}$ and ${\gamma_{j}^{t}}$ respectively, their independence is clear. The independence is less trivial in the $SO(D-1)$ scalar sector involving $\{I_{W},I,h\}$. At first sight we might associate those three scalars with $\{\rho,\Lambda,\psi\}$ respectively, however when we get $\rho(I_{W},I,\psi,\Gamma)$ from $I_{W}$ and $I$ and plug back in the trace $h$, all helicities will appear under time derivative. In order to avoid it we must first get rid of $\ddot{\psi}$ inside $I_{W}$ by redefining $\Gamma\equiv\bar{\Gamma}+\dot{\psi}/2$. So we end up with a change of variables from $\{\rho,\Lambda,\psi,\Gamma\}$ to $\{I_{W},I,h,\bar{\Gamma}\}$. We can check that the Jacobian does not depend upon time derivatives, namely,
@@
We can show as well that the $SO(D-1)$ scalar sector of $I_{A}^{WTD}$ (i.e. $\{I_{W},I_{WT},h\}$) form an independent basis. It turns out that $\det[\mathbb{M}^{(W)}_{TD}]=\det[\mathbb{M}_{TD}]=(D-2)^{2}\nabla^{4}$, where $M^{(W)}_{TD}$ is the matrix appearing in the change of coordinates from $\{\rho,\Lambda,\psi,\Gamma\}$ to $\{I_{W},I_{WT},h,\bar{\Gamma}\}$, where $I_{WT}=(D-2)(\rho+2\nabla^{2}\Lambda-\nabla^{2}\psi)/D$.
@@ 3.1 NSTDiff spectrum via a shortcut
In the present section we extend the theories presented in the last section by replacing the symmetric rank-2 fundamental field by a general rank-2 tensor with both symmetric $(h_{\mu\nu}=h_{\nu\mu})$ and antisymmetric ($B_{\mu\nu}=-B_{\nu\mu}$) parts: $e_{\mu\nu}=h_{\mu\nu}+B_{\mu\nu}$. Starting from a general theory of second order in derivatives and requiring symmetry under nonsymmetric TDiff (NSTDiff) : $\delta e_{\mu\nu}=\partial_{\nu}\,\epsilon_{\mu}^{T}$ we end up with two possibilities. They both depend in principle on three arbitrary real parameters [^7] We could have defined $\delta e^{T}_{\mu\nu}=a\,\partial_{\mu}\epsilon^{T}_{\nu}+(1-a)\partial_{\nu}\epsilon^{T}_{\mu}$ with some arbitrary $a$, however this could be undone by redefining the fundamental field taking combinations of $e^{T}_{\mu\nu}$ and $e^{T}_{\nu\mu}$.. In the first case $B_{\mu\nu}$ is decoupled from $h_{\mu\nu}$,
~$\tilde{{\cal L}}_{NST}(a,b,\tilde{c})=\frac{1}{2}h^{\alpha\beta}\,\Box\,h_{\alpha\beta}+(\partial^{\mu}h_{\mu\beta})^{2}+a\,h\,\partial^{\mu}\partial^{\nu}h_{\mu\nu}-\frac{b}{2}h\,\Box h+\,\tilde{c}\,[B_{\mu\nu}\,\frac{\Box}{2}\,B^{\mu\nu}+(\partial^{\mu}B_{\mu\nu})^{2}].$ (24)
where $\tilde{c}$ is a real constant. The model (24) is invariant under
@@
The question arises: what is the particle content of ${\cal L}_{NST}(a,b,c)$ We follow now the same shortcut of the symmetric case. We search for an r-shift $e_{\mu\nu}\to e_{\mu\nu}+r\,\eta_{\mu\nu}\,h$ which might split ${\cal L}_{NST}(a,b,c)$ into two pieces whose content could be investigated separately. First, assuming that $a\neq a_{D}(c)$, we can always choose $r=(2c+1-a)/[a\,D-2(1+c)]$ which leads to a sum of a general NSDiff model plus a kinetic term for the trace $h$. Namely,
@@
where the constant $B$ is an analytic function of $(a,b,c)$ except at $a=a_{D}(c)$ whose details are not important for us as we explain below. The model ${\cal L}_{NSD}(c)$ defined in (31) is invariant under (28). Regarding the case $a=a_{D}(c)=2(1+c)/D$, we can, similarly to the symmetric case, split ${\cal L}[a_{D}(c),b,c]$ into a NSWTDiff model plus a kinetic term for the trace without need of an r-shift,
${\cal L}_{NST}[a_{D}(c),b,c]={\cal L}[a_{D}(c),b_{D}(c),c]+\frac{\tilde{B}}{2}\,h\,\Box\,h\equiv{\cal L}_{NSWTD}(c)+\frac{\tilde{B}}{2}\,h\,\Box\,h\quad,$ (32)
@@
We will show in the next subsection that ${\cal L}_{NSD}(c)$ can be written in terms of NSDiff gauge invariants $I_{A}^{NSD}$, $A=1,2,\cdots,N_{I}^{NSD}$ where $N_{I}^{NSD}=D^{2}-[D+D(D-1)/2-(D-1)]$. The transverse condition $\partial^{\mu}\epsilon_{\mu}^{T}=0$ reduces one unit of gauge parameters which leads to one extra gauge invariant which we chose to be the trace $h$ again. Thus, we will use the following basis for the NSTDiff invariants $I_{\tilde{A}}^{NSTD}=\left\{I_{A}^{NSD},h\right\}$. In order to find out the content of (31) we only need to know the content of ${\cal L}_{NSD}(c)$ which, for arbitrary values of $c$, has first appeared in [25] where its content has been investigated [^8] The model ${\cal L}_{NSD}(c)$ corresponds to the ${\cal L}(a_{1})$ theory in [25] with the map $c=a_{1}-1/4$. via a Lagrangian method based on the introduction of an auxiliary vector field and confirmed in [26] via the usual hamiltonian Dirac method. The model ${\cal L}_{NSD}(c)$ has been rediscovered recently in [27], where it is named type 3, as a special case of linearized NGR about Minkowski space. In the next subsection, we show via gauge invariants that it describes, see also [25], a physical massless spin-2 particle and a spin-0 particle which is physical if either $c>0$ or $c<-c_{D}$ where
@@
Regarding the special case $a=a_{D}(c)$ in (32), we have exactly the same content of (31) under the same conditions, i.e., a physical massless SST theory whenever $B>0$ and either $c>0$ or $c<-c_{D}$. The coincidence stems from the fact that, as in the symmetric case of the last section, we have ${\cal L}_{NSWTD}(h_{\mu\nu})={\cal L}_{NSD}[\bar{h}_{\mu\nu}(h)]$, likewise for the invariants $I_{A}^{NSWTD}(h_{\mu\nu})=I_{A}^{NSD}[\bar{h}_{\mu\nu}(h)]$. Since they are all independent quantities as we will show here, the same spectrum must hold under the same conditions.

### v2

# Linearized transverse diffeomorphism invariant spin-2 theories via gauge invariants

@@ Abstract
We analyze the particle spectrum of a second-order (in derivatives) theory based on a rank-2 tensor field with both symmetric and antisymmetric components. By demanding the existence of a propagating massless spin-2 particle and invariance under linearized transverse diffeomorphisms, we derive a new class of stable models with two massless scalars and a single massless spin-2 particle. A natural nonlinear completion is proposed in terms of a dynamical metric field. The identification of the spectrum is carried out using a fully Lagrangian, gauge-invariant approach which makes use of Bardeen variables in a constructive manner. The approach significantly reduces the number of steps in the spectrum determination in some cases.
@@ 1 Introduction
The TDiff models have some general interesting aspects. Notably, they have an integration constant with the same mass dimension as the cosmological constant which may have interesting phenomenological consequences in scale invariant theories of gravity [^4] In particular, in TDiff theories there naturally appears an integration constant with the same mass dimension of the cosmological constant which may have relevant application to the dark energy problem, see [8]. [8]. In WTDiff gravity, the constant plays the role of a cosmological constant that does not receive quantum corrections, see discussion in [7]. It is also worth commenting that TDiff models (with or without Weyl symmetry) cannot become massive in their spin-2 sector [5]. Given the stringent experimental bounds on the graviton mass ($m_{g}\lesssim 10^{-23}~\mathrm{eV}$), the TDiff symmetry seems to be more natural than the usual Diff one (Einstein-Hilbert) which can become massive at linearized level [9] and beyond [10, 11], see also [12].
@@
The aim of the present work is twofold, first we merge the previous mentioned directions of modifying gravity at linearized level and investigate the most general Lorentz invariant quadratic theory of second order in derivatives invariant under TDiff and described by a full rank-2 tensor. Thus, paving the way for non linear completions. Secondly, the content of the model is examined by using an explicitly gauge invariant approach based on an improved use of Bardeen variables [19] which turns out to be a very convenient way of unravelling the particle content of quadratic tensor models.
@@ 2.1 TDiff spectrum via a shortcut
The most general massless TDiff invariant local spin-2 model of second order in derivatives, described only in terms of a symmetric rank-2 field, is given by the following Lagrangian:
@@
The first two terms are required for massless spin-2 propagation without spin-1 ghosts [5]. The last two terms depend on two real constants: $(a,b)$. Up to surface terms, neglected throughout this work, the lagrangian in (1) is TDiff invariant [^5] Here we use the index “TT” for space-time ($\partial^{\mu}v_{\mu}^{T}$=0) and “tt” for space ($\partial_{j}v_{j}^{t}=0$) transverse quantities. ($\partial^{\mu}\epsilon^{T}_{\mu}=0$) for arbitrary $(a,b)$,
@@
In the first case, we have the linearized Einstein-Hilbert [^6] More precisely, there is a continuous family of models $(a(r),b(r))$ equivalent to the point $(a,b)=(1,1)$ via what we call r-shifts: $h_{\mu\nu}\to h_{\mu\nu}+r\,\eta_{\mu\nu}\,h$ for which the enhancement still holds, though (3) is modified accordingly, while the case $(a,b)=(a_{D},b_{D})$ is a fixed point of the r-shifts [5]. theory $\mathcal{L}_{TD}(1,1)=\mathcal{L}_{EH}$ where TDiff is enhanced to Diff. In the second one, the symmetry becomes Weyl + TDiff (WTDiff). In both cases, the model describes only massless spin-2 particles in $D=4$ with only two degrees of freedom corresponding to the helicities $\pm 2$.
For an otherwise arbitrary couple $(a,b)$, a quite lengthy analysis, see [5],[20] and [21], reveals that we have in general (D1)(D2)/2(D-1)(D-2)/2 degrees of freedom corresponding to a massless spin-2 and a massless spin-0 particle in $D=4$. A key quantity is $f_{D}(a,b)$:
@@
where $A,B=0,\cdots,N_{I}$. The matrix ${\cal M}_{AB}$ depends on space and time derivatives. The invariant indices $A,B$ are just labeling indices. They are deprived of meaning regarding Lorentz transformations. Regarding their range, $N_{I}$ it is the number of independent gauge invariants. We will show that $N_{I}$ equals the number of fields minus the number of independent gauge parameters, generically,
@@
When we go from Diff to TDiff we have one less independent gauge parameter due to $\partial^{\mu}\epsilon_{\mu}^{T}=0$, consequently, from (7) we must have one extra gauge invariant. Since any Diff invariant must be also TDiff invariant and the trace $h$ is TDiff invariant by itself, we can consider $I_{\tilde{A}}^{TDiff}=\left\{I_{A}^{Diff},h\right\}$ as a basis of independent TDiff invariants. The trace $h$ cannot be a function of the Diff invariants $I_{A}^{Diff}$, otherwise it would be also Diff invariant which is certainly not the case, recall $\delta\,h=2\,\partial\cdot\epsilon\neq 0$. So if we were able to rewrite (1) as a sum of the linearized EH theory and a term that depends only on the trace $h$ we could immediately identify the particle content of (1). Indeed, there is a unique field redefinition which accomplishes it, namely, after $h_{\mu\nu}\to h_{\mu\nu}+r\,\eta_{\mu\nu}h$ with r=(1a)/(aD2)r=(1-a)/(aD-2), assuming $a\neq 2/D$, the TDiff theory (1) can be written, in terms of the new field, as
@@
Thus, although the field $h_{\mu\nu}$ in ${\cal L}_{EH}(h_{\mu\nu})$ is not traceless, it turns out that the particle content of ${\cal L}_{TD}(a,b)$ can be read off from the two independent terms in (8), in full agreement with the more laborious analysis of [5], [20] and [21]. Namely, it corresponds to a physical massless spin-2 particle plus a massless scalar particle which is physical (ghostly) if $f_{D}(a,b)>0$ ($f_{D}(a,b)<0$) and disappears at $f_{D}(a,b)=0$.
@@
where ${\cal L}_{WTD}(h_{\mu\nu})={\cal L}_{TD}(a_{D},b_{D})$, with $f_{D}(a_{D},b)=(D-2)(b_{D}-b)$ and $I_{A}^{WTD}$ are the invariants under WTDiff. Notice that $N_{I}^{WTD}=N_{I}^{Diff}$ since the longitudinal component of $\epsilon_{\mu}$ is replaced by the Weyl gauge parameter such that the number of independent gauge parameters remains the same. On the other hand, as we go from WTDiff to TDiff, the broken Weyl symmetry gives rise to one extra invariant which may be again the trace $h$ that now breaks the Weyl symmetry such that we can choose $I_{\tilde{A}}^{TDiff}=\left\{I_{A}^{WTD},h\right\}$ as another basis for the TDiff invariants. The physical content of ${\cal L}_{TD}(a_{D},b)$ factorizes into the content of the two terms in (9), once again in agreement with [5], [20] and [21].
@@ 2.2 Diff, WTDiff and TDiff gauge invariants
we can express $\epsilon_{\mu}=\epsilon_{\mu}(\delta h_{\mu\nu})$. It is convenient to decompose the gauge parameters in terms of SO(D1)SO(D-1) scalars and vectors,
@@
where $\omega_{ij}=\partial_{i}\partial_{j}/\nabla^{2}$ and $\theta_{ij}=\delta_{ij}-\omega_{ij}$ are longitudinal and transverse space projection operators. Using (12) into (10), the remaining D(D+1)/2DD(D+1)/2-D equations become invariance equations $\delta I_{A}^{Diff}=0$, where $I_{A}^{Diff}=\{I_{ij}^{t},I_{j}^{t},I_{00}\}$ are the Diff invariants:
@@
They represent a total of $N_{I}^{Diff}=1+(D-2)+[(D-1)D/2-(D-1)]=(D-1)D/2$ independent Diff invariants in agreement with (7): $N(h_{\mu\nu})-N(\epsilon_{\mu})=D(D+1)/2-D$. It is convenient to redefine them on physical grounds. Since the two helicity states $\pm 2$ in $D=4$ correspond to the traceless and transverse components of $h_{ij}$, we decompose $I_{ij}^{t}$ into its traceless ($I_{ij}^{tt}$) and trace ($I$) components. Respectively,
@@
$I=(D-2)\nabla^{2}\Lambda\ ;\qquad h=-\rho+\nabla^{2}(D-2)\Lambda+\nabla^{2}\psi\$(21)
We can always associate $I_{ij}^{tt}$ and $I_{j}^{t}$ with the field components $\Psi_{ij}^{tt}$ and $\gamma_{j}^{t}$} respectively, their independence is clear. The independence is less trivial in the SO(D1)SO(D-1) scalar sector involving $\{I_{W},I,h\}$. At first sight we might associate those three scalars with $\{\rho,\Lambda,\psi\}$ respectively, however when we get $\rho(I_{W},I,\psi,\Gamma)$ from $I_{W}$ and $I$ and plug back in the trace $h$, all helicities will appear under time derivative. In order to avoid it we must first get rid of $\ddot{\psi}$ inside $I_{W}$ by redefining $\Gamma\equiv\bar{\Gamma}+\dot{\psi}/2$. So we end up with a change of variables from $\{\rho,\Lambda,\psi,\Gamma\}$ to $\{I_{W},I,h,\bar{\Gamma}\}$. We can check that the Jacobian does not depend upon time derivatives, namely,
@@
We can show as well that the SO(D1)SO(D-1) scalar sector of $I_{A}^{WTD}$ (i.e. $\{I_{W},I_{WT},h\}$) form an independent basis. It turns out that $\det[\mathbb{M}^{(W)}_{TD}]=\det[\mathbb{M}_{TD}]=(D-2)^{2}\nabla^{4}$, where $M^{(W)}_{TD}$ is the matrix appearing in the change of coordinates from $\{\rho,\Lambda,\psi,\Gamma\}$ to $\{I_{W},I_{WT},h,\bar{\Gamma}\}$, where $I_{WT}=(D-2)(\rho+2\nabla^{2}\Lambda-\nabla^{2}\psi)/D$.
@@ 3.1 NSTDiff spectrum via a shortcut
In the present section we extend the theories presented in the last section by replacing the symmetric rank-2 fundamental field by a general rank-2 tensor with both symmetric $(h_{\mu\nu}=h_{\nu\mu})$ and antisymmetric ($B_{\mu\nu}=-B_{\nu\mu}$) parts: $e_{\mu\nu}=h_{\mu\nu}+B_{\mu\nu}$. Starting from a general theory of second order in derivatives and requiring symmetry under nonsymmetric TDiff (NSTDiff): $\delta e_{\mu\nu}=\partial_{\nu}\,\epsilon_{\mu}^{T}$ we end up with two possibilities. They both depend in principle on three arbitrary real parameters [^8] We could have defined $\delta e_{\mu\nu}=a\,\partial_{\mu}\epsilon^{T}_{\nu}+(1-a)\partial_{\nu}\epsilon^{T}_{\mu}$ with some arbitrary $a$, however this could be undone by redefining the fundamental field taking combinations of $e_{\mu\nu}$ and $e_{\nu\mu}$..
In the first case $B_{\mu\nu}$ is decoupled from $h_{\mu\nu}$,
$\tilde{{\cal L}}_{NST}(a,b,\tilde{c})=\frac{1}{2}h^{\alpha\beta}\,\Box\,h_{\alpha\beta}+(\partial^{\mu}h_{\mu\beta})^{2}+a\,h\,\partial^{\mu}\partial^{\nu}h_{\mu\nu}-\frac{b}{2}h\,\Box h+\,\tilde{c}\,[B_{\mu\nu}\,\frac{\Box}{2}\,B^{\mu\nu}+(\partial^{\mu}B_{\mu\nu})^{2}].$ (24)
where $\tilde{c}$ is a real constant. The new model (24) is invariant under
@@
The question arises: what is the particle content of ${\cal L}_{NST}(a,b,c)$ ? We follow now the same shortcut of the symmetric case. We search for an r-shift $e_{\mu\nu}\to e_{\mu\nu}+r\,\eta_{\mu\nu}\,h$ which might split ${\cal L}_{NST}(a,b,c)$ into two pieces whose content could be investigated separately. First, assuming that $a\neq a_{D}(c)$, we can always choose r=(2c+1a)/[aD2(1+c)]r=(2c+1-a)/[a\,D-2(1+c)] which leads to a sum of a general NSDiff model plus a kinetic term for the trace $h$. Namely,
@@
where the constant $B$ is an analytic function of $(a,b,c)$, except at $a=a_{D}(c)$, whose details are not important for us as we explain below. The model ${\cal L}_{NSD}(c)$ defined in (31) is invariant under (28). It has first appeared [^9] The model ${\cal L}_{NSD}(c)$ corresponds to the ${\cal L}(a_{1})$ theory in [25] with the map $c=a_{1}-1/4$. in [25] where one has searched for the most general Lorentz invariant massive spin-2 model of second order in derivatives of an arbitrary rank-2 tensor. Those requirements have led to a family of massive models given in terms of an arbitrary real parameter, see [25], which includes the paradigmatic Fierz-Pauli model [9]. Its massless limit is given by ${\cal L}_{NSD}(c)$.
Regarding the case $a=a_{D}(c)=2(1+c)/D$, we can, similarly to the symmetric case, split ${\cal L}_{NST}[a_{D}(c),b,c]$ into a NSWTDiff model plus a kinetic term for the trace without need of an r-shift,
${\cal L}_{NST}[a_{D}(c),b,c]={\cal L}_{NST}[a_{D}(c),b_{D}(c),c]+\frac{\tilde{B}}{2}\,h\,\Box\,h\equiv{\cal L}_{NSWTD}(c)+\frac{\tilde{B}}{2}\,h\,\Box\,h\quad,$ (32)
@@
We will show in the next subsection that ${\cal L}_{NSD}(c)$ can be written in terms of NSDiff gauge invariants $I_{A}^{NSD}$, $A=1,2,\cdots,N_{I}^{NSD}$ where $N_{I}^{NSD}=D^{2}-[D+D(D-1)/2-(D-1)]$. The transverse condition $\partial^{\mu}\epsilon_{\mu}^{T}=0$ reduces one unit of gauge parameters which leads to one extra gauge invariant which we chose to be the trace $h$ again. Thus, we will use the following basis for the NSTDiff invariants $I_{\tilde{A}}^{NSTD}=\left\{I_{A}^{NSD},h\right\}$. In order to find out the content of (31) we only need to know the content of ${\cal L}_{NSD}(c)$. Its content has been investigated via a Lagrangian method based on the introduction of an auxiliary vector field and confirmed in [26] via the usual hamiltonian Dirac method. The model ${\cal L}_{NSD}(c)$ has been rediscovered recently in [27], where it is named type 3, as a special case of linearized NGR about Minkowski space. In the next subsection, we show via gauge invariants that it describes, see also [25], a physical massless spin-2 particle and a spin-0 particle which is physical if either $c>0$ or $c<-c_{D}$ where
@@
Regarding the special case $a=a_{D}(c)$ in (32), we have exactly the same content of (31) under the same conditions, i.e., a physical massless SST theory whenever $B>0$ and either $c>0$ or $c<-c_{D}$. The coincidence stems from the fact that, as in the symmetric case of the last section, we have ${\cal L}_{NSWTD}(h_{\mu\nu})={\cal L}_{NSD}[\bar{h}_{\mu\nu}(h)]$, likewise for the invariants $I_{A}^{NSWTD}(h_{\mu\nu})=I_{A}^{NSD}[\bar{h}_{\mu\nu}(h)]$. Since they are all independent quantities as we will show here, the same spectrum must hold under the same conditions. In table 1 we summarize the linearized new models suggested here.
Model Formula Symmetries Restrictions
[…]
$\delta B_{\mu\nu}=\partial_{\nu}\,\lambda^{T}_{\mu}-\partial_{\mu}\,\lambda^{T}_{\nu}$
${\cal L}_{NST}(c,B)$    (31) $\delta e_{\mu\nu}=\partial_{\nu}\,\epsilon_{\mu}^{T}+\Lambda_{\mu\nu}^{T}$ $c>0$   or   $c<-c_{D}$ and $B>0$
${\cal L}_{NST}(a_{D},b,c)$ (32)   and   (26) $\delta e_{\mu\nu}=\partial_{\nu}\,\epsilon_{\mu}^{T}+\Lambda_{\mu\nu}^{T}$ $c>0$   or   $c<-c_{D}$ and $\tilde{B}>0$
Table 1: New models with their local symmetries and restrictions on the parameter’s space. They all have the same particle content, one spin-2 and two spin-0 massless modes (all physical). Notice that $(h_{\mu\nu},B_{\mu\nu})=(e_{(\mu\nu)},e_{[\mu\nu]})$